OpenStudy (anonymous):

GIVING AWAY A MEDAL: Use Euler’s Formula to find the missing number. Faces: 19 Vertices: 11 Edges: ? A) 29 B) 28 C) 27 D) 31

OpenStudy (anonymous):

@iGreen Do you know what this one is?

OpenStudy (igreen):

No. I never heard of Euler's Formula. Sorry.

OpenStudy (anonymous):

Thanks anyways.

OpenStudy (mathmath333):

\(\large euler's ~~formula~~ (f+v-e=2)\)

OpenStudy (igreen):

Oh, well.. Given what @mathmath333 said: I'm betting f = faces, v = vertices, and e = edges.

OpenStudy (anonymous):

Yes.

OpenStudy (igreen):

Plug in what you know: 19 + 11 - e = 2 Add: 30 - e = 2 Subtract 30 to both sides: -e = 32 Multiply -1 to both sides: e = -32

OpenStudy (igreen):

Okay, I did something wrong..nevermind. xD

OpenStudy (anonymous):

Lol, thats okay. You tried. I have absolutely no idea how to do these.

OpenStudy (mathmath333):

sry 28 lol

OpenStudy (anonymous):

Thank you.